# Divide, Ensemble and Conquer: The Last Mile on Unsupervised Domain Adaptation for On-Board Semantic Segmentation

### v1

@@ I Introduction
In this paper, we propose a novel multi-source framework called Divide, Ensemble and Conquer (DEC). It follows a divide-and-conquer strategy to train semantic segmentation models and fuse their predictions to generate the semantic segmentation mask. In our first step, we apply the division strategy to train several models on categorized semantic classes, called category models, similarly to how humans semantically label an RGB image using AI tools [20, 21] (first background, then large objects, and finally small ones). The categorization simplifies the number of classes learnt by each category model, expecting a better performance and fast convergence. The next step involves the training of ensemble model, a DeepLabv3+ [22] network, which learns to fuse segmentation masks generated by each category model into one mask. The framework is illustrated in Fig. 1. DEC exhibits compatibility with a wide range of UDA methods, providing consistent improvement. To ensure a fair comparison, we reproduce the results of previous UDA methods with the multi-source dataset. DEC achieves state-of-the-art performance on Cityscapes, BDD100K [23], and Mapillary Vistas [24], resulting in mIoU of 78.7, 62.4, and 74.8, respectively. We enhance the existing state-of-the-art reported in [11] by 1.2, 2.1, and 0.6 points, respectively. Our performance is only 2.8 and 4.7 points behind SL for Cityscapes and Mapillary Vistas, surpassing the SL for BDD by 0.9 points. Overall, DEC propels progress in the challenging last mile of UDA through a simple yet effective divide-and-conquer strategy and ensemble method.
@@ II-B UDA for Semantic Segmentation
Most UDA works relies on single-source synthetic dataset [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 9, 10, 11, 46, 47, 48, 34]. Hoyer et al. [9] successfully use the novel vision-transformers [49] to solve UDA for semantic segmentation. They employ a teacher-student self-training framework with additional training strategies, such as computing ImageNet feature distances. Furthermore, several new works propose techniques to improve performance on top of existing frameworks (DACS, Daformer, etc.) Hoyer et al. [10, 11] adopt a multi-resolution image cropping approach to capture context dependencies and propose a masked image consistency module to learn spatial context relations of the target domain. Chen et al. [12] explore the pixel-to-pixel and patch-to-patch relation for regularizing the segmentation feature space. These works make significant progress for GTA5 $\rightarrow$ Cityscapes but always need to design specific strategies (e.g., RCS and FD) for class unbalance due to the drawback of GTA5. These specific strategies do not generalize effectively to multi-source datasets (see Section IV-E1). On the other hand, [9, 10, 11, 12] add extra pipelines for training on top of previous works, which makes the UDA become more and more complex and hard to train, e.g., [11] is on the top of DACS[8], Daformer[9] and HRDA[10]. Our DEC method is entirely data-driven and does not require an extra algorithm for UDA training.
@@ III Method
This section details the proposed framework, DEC, grounded in the divide-and-conquer strategy. The framework comprises two components: category models and an ensemble model. Section. III-A discusses the division strategy and category models, which segment images into category masks. Subsequently, Section. III-B explains the ensemble model, designed to fuse predictions derived from category models.
@@ III-A Division Strategy
Employing the divide-and-conquer strategy, we deconstruct the semantic segmentation task into subtasks that group classes into categories and train category models to segment different classes. In particular, we group $N_{C}$ classes into $N_{G}$ categories guided by specific criteria which encompasses various factors, such as how humans semantically label an RGB image by AI tools[20, 21] (e.g. commencing with the background, progressing to large objects, and concluding with smaller ones). To account for non-category classes (classes that do not belong to a particular category), we introduce a class named other category to denote them. It means a specific class is assigned only in one category label for a given pixel, while other category labels maintain the designation as other category. This class helps category models learn features of non-category classes in training, improving the accuracy of classes that belong to a category. Subsequently, we remap source labels $\mathcal{Y^{S}}$ to source category labels $\mathcal{Y}^{S}_{j}$ based on the defined $N_{G}$ categories where the class will be retained in the category label if it belongs to this category. The procedure of generating category labels is shown in Algorithm 1. An example of remapping is shown in Fig. 3.
@@
After generating source category labels, we train target category models, denoted as $\{f_{\theta^{j}}^{T}\}_{j=1}^{N_{G}}$, using existing UDA methods. Each target category model $f_{\theta^{j}}^{T}$ is then employed to segment target images $\mathcal{X}^{T}$ into target category masks $\hat{\mathcal{Y}}^{T}_{j}$, encompassing $N_{j}$ classes.
III-B Ensemble Model
@@
By division strategy detailed in Section. III-A, target images $\mathcal{X}^{T}$ are segmented into target category masks $\hat{\mathcal{Y}}^{T}_{j}$. Conflicts arise when different classes are predicted for the same pixel value among the target category masks. Note that popular ensemble methods like majority voting and averaging are not suitable to solve these conflicts. Due to the aforementioned division strategy, category models do not share classes, thus ensemble methods that depends on a consensus between models can not be applied. Hence, we introduce an ensemble model $E_{\theta}$ designed to extract features from each target category mask $\hat{\mathcal{Y}}^{T}_{j}$ and fuse them into the final semantic segmentation mask $\hat{\mathcal{Y}}^{T}$.
@@
Training Step. The training step is implemented in Algorithm 2 and illustrated in Fig. 2. In the training dataset ($(\{\hat{\mathcal{Y}}^{S}_{j}\}_{j=1}^{N_{G}},\mathcal{Y}^{S})$, there are $N_{G}$ category masks for one source label. To enable the model to process this kind of input, we stack them (similar to how we treat RGB images) into a pseudo-image $\mathcal{X}^{E}$ where each channel corresponds to a source category mask $\hat{\mathcal{Y}}^{S}_{j}$. The $\mathcal{X}^{E}$ channels follow the same order as the $N_{G}$ categories defined. We use pixel-wise cross-entropy as the loss function $\mathcal{L}^{E}$ to train the ensemble model like semantic segmentation task
@@ IV-B2 Source Category Models
Source category models are employed to train the ensemble model. The network architecture uses DeepLabV3+ [55] with ResNet101 [56] as the backbone, loaded with ImageNet pretrained weights. Training adopts SGD [57] with a learning rate of $2\times 10^{-3}$, linear learning rate warmup over 1k iterations, and polynomial decay. The model undergoes training on a batch of eight images with a crop size 1024$\times$512 for 90k iterations.
@@ IV-E Ablation Study
TABLE IV: Performance of random and other feasible division strategies on Musketeers $\rightarrow$ Cityscapes.
Random Division
Other Feasible Division
IV-E1 Source data
TABLE V: Rare Class Sampling (RCS) and Thing-class ImageNet Feature Distance (FD) are applied to a multi-source dataset, with the results averaged over three random seeds.
@@ IV-E1 Source data
DEC use multi-source datasets to train category and ensemble models. For category models, as stated in Section I, strategies tailored for single-source synthetic dataset are often susceptible to overfitting and instability. MIC and HRDA exhibit noteworthy advancements in GTA $\rightarrow$ Cityscapes based on RCS and FD while these strategies demonstrate suboptimal performance in multi-source scenarios (see Table V). On the other hand, the ensemble model needs enough training data since there is no appropriate pre-trained backbone. The variability of single-source dataset is insufficient for developing a robust ensemble model. Table VI presents the mIoU of ensemble models trained on single-source and multi-source datasets. The results demonstrate that the ensemble model trained on multi-source datasets effectively fuses category masks into the final mask, e.g., the Human/Cycle category (see Fig. 5 (c)) correctly segments the dotted area, whereas the ensemble model trained with GTA5 fuse dotted area incorrectly (see Fig. 5 (g)). Thus, it is necessary to train the ensemble model with multi-source datasets to obtain robust and reliable final masks.
TABLE VI: The influence of single-source and multi-source training datasets on the ensemble model. We show the performance of GTA5 $\rightarrow$ Cityscapes and Musketeers $\rightarrow$ Cityscapes on the top of MIC.
@@ IV-E2 Division Strategy
DEC groups classes into categories to decrease the complexity of the model and improve segmentation. However, DEC requires a feasible division strategy (including but not limited to Table I) to maintain the contextual information between different classes. Choosing proper categories is not an arbitrary task. We demonstrate in Table VII how randomly pick some classes as a group to generate some categories. It is seen that random division does not push the performance for Musketeers $\rightarrow$ Cityscapes in Table IV. Additionally, we show the performance of other feasible division strategies such as dividing classes in two categories and three categories in Table IV. These division strategies are generated by combining categories in Table I, e.g., BV denotes a category containing classes from Background and Vehicle; BV+HT means this division strategy splits all classes into two groups. These division strategies achieve a gain of at least $+0.8$ for Musketeers $\rightarrow$ Cityscapes compared to MIC. Fig. 6 provides a qualitative result where these feasible strategies can effectively improve segmentation, e.g., sidewalk.
TABLE VII: Random division strategy. Each row in the second column is a category.
@@ IV-E3 Ensemble Model
Backbone mIoU (Cityscapes) mIoU (BDD100K) mIoU (Mapillary Vista)
@@
Backbone: Table IX illustrates the performance of the ensemble model with various backbones for Musketeers $\rightarrow$ Cityscapes, BDD100K and Mapillary Vista. While ResNet18 attains optimal performance, the disparities among different backbones are relatively minor. Given the trade-off between model complexity and training time, we select ResNet18 as the preferred backbone for the ensemble model.
TABLE IX: The performance of ensemble model with different backbones for three target domains.
Backbone mIoU (Cityscapes) mIoU (BDD100K) mIoU (Mapillary Vista)
ResNet18 79.1 63.2 75.0
ResNet34 78.9 63.1 74.7
ResNet50 78.2 62.6 74.2
ResNet101 78.9 63.1 75.0
Optimizer: Tab. X shows the SGD and AdamW with different learning rates to train the ensemble model for Musketeers $\rightarrow$ Cityscapes. It can be seen that AdamW is more compatible with the ensemble model than SGD.
TABLE X: The performance of SGD and AdamW optimiser with different learning rates for Musketeers $\rightarrow$ Cityscapes.
@@
EMA: Tab. XI shows the influence of different EMA coefficients $\alpha$ for Musketeers $\rightarrow$ Cityscapes, BDD100K and Mapillary Vista. It is observed that a larger value of $\alpha$ enhances the generalisation of the ensemble model, leading to optimal performance for each target dataset.
TABLE XI: The impact of EMA coefficient $\alpha$ for three target domains.
EMA Coefficient $\alpha$ mIoU (Cityscapes) mIoU (BDD100K) mIoU (Mapillary Vista)
@@
IV-F The Last Mile to SL
@@
Some classes still exhibit gaps compared to SL. For example, the sidewalk, an important class on autonomous driving, shows a 5 points difference from SL on Musketeers $\rightarrow$ Cityscapes. Fig. 7 shows some qualitative results for the sidewalk in both DEC and SL. The error in DEC is primarily found at the bottom of vehicles, where the sidewalk is misclassified as road. We expect these errors to have an insignificant influence on autonomous driving, as the vehicles in these areas are classified correctly. Moreover, since the ground truth is manually labelled, human bias is inevitably introduced into the ground truth. This is why some classes show discrepancies compared to SL, such as terrain, fence, and pole. Fig. 8 illustrates these human biases in Cityscapes. Furthermore, there are also human bias in Mapillary Vistas. Fig. 9 indicates annotation bias among terrain, building, sky, and vegetation. These classes are labelled differently between synthetic and real datasets. This explains the large gap (e.g., vegetation, terrain) in our method for Mapillary Vista.

### v2

# Divide, Ensemble and Conquer: The Last Mile on Unsupervised Domain Adaptation for Semantic Segmentation

@@ I Introduction
In this paper, we propose a novel multi-source framework called Divide, Ensemble and Conquer (DEC). It follows a divide-and-conquer strategy to train semantic segmentation models and fuse their predictions to generate the semantic segmentation mask. In our first step, we apply the division strategy to train several models on categorized semantic classes, called category models, similarly to how humans semantically label an RGB image using AI tools [20, 21] (first background, then large objects, and finally small ones). The categorization simplifies the number of classes learnt by each category model, expecting a better performance and fast convergence. The next step involves the training of an ensemble model, a DeepLabv3+ [22] network, which learns to fuse segmentation masks generated by each category model into one mask. The framework is illustrated in Fig. 1. DEC exhibits compatibility with a wide range of UDA methods, providing consistent improvement. To ensure a fair comparison, we reproduce the results of previous UDA methods with the multi-source dataset. DEC achieves state-of-the-art performance on Cityscapes, BDD100K [23], and Mapillary Vistas [24], resulting in mIoU of 78.7, 62.4, and 74.8, respectively. We enhance the existing state-of-the-art reported in [11] by 1.2, 2.1, and 0.6 points, respectively. Our performance is only 2.8 and 4.7 points behind SL for Cityscapes and Mapillary Vistas, surpassing the SL for BDD by 0.9 points. Overall, the main contributions of our work are as follows:
  1. 1.
    A novel division approach: DEC categorizes semantic classes into groups (e.g., background, large objects, and small objects) and trains category models. This approach simplifies the feature space for each category model, leading to improved UDA performance.
  2. 2.
    An ensemble-based fusion mechanism: DEC leverages an ensemble model to fuse the outputs from category models into the final segmentation mask. This fusion approach ensures robustness and compatibility with various UDA methods, providing consistent improvements across datasets.
  3. 3.
    State-of-the-art performance on multiple benchmarks: DEC achieves competitive results on Cityscapes, BDD100K, and Mapillary Vistas, surpassing existing UDA methods and narrowing the gap with SL.
  4. 4.
    Flexibility and ease of use: Besides the label division strategy, DEC does not rely on specialized strategy or additional training parameters. It can be easily integrated with various UDA methods without changing the baseline model architecture.
@@ II-A Ensemble Learning
Ensemble learning enhances the generalization performance of deep learning models by combining the predictions of multiple models, reducing variance and mitigating overfitting. This approach enables us to decompose the semantic segmentation task into smaller, more manageable subtasks, each focused on a specific set of semantic categories. By training specialized models for each category, we can leverage the strengths of individual models (category experts), resulting in more robust and accurate segmentation result.
@@ II-B UDA for Semantic Segmentation
Most UDA works relies on single-source synthetic dataset [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 9, 10, 11, 46, 47, 48, 34]. Hoyer et al. [9] successfully use the novel vision-transformers [49] to solve UDA for semantic segmentation. They employ a teacher-student self-training framework with additional training strategies, such as computing ImageNet feature distances. Furthermore, several new works propose techniques to improve performance on top of existing frameworks (DACS, Daformer, etc.) Hoyer et al. [10, 11] adopt a multi-resolution image cropping approach to capture context dependencies and propose a masked image consistency module to learn spatial context relations of the target domain. Chen et al. [12] explore the pixel-to-pixel and patch-to-patch relation for regularizing the segmentation feature space. These works make significant progress for GTA5 $\rightarrow$ Cityscapes but always need to design specific strategies (e.g., RCS and FD) for class unbalance due to the drawback of GTA5. These specific strategies do not generalize effectively to multi-source datasets (see Section 5). On the other hand, [9, 10, 11, 12] add extra pipelines for training on top of previous works, which makes the UDA become more and more complex and hard to train, e.g., [11] is on the top of DACS[8], Daformer[9] and HRDA[10]. Our DEC method is entirely data-driven and does not require an extra algorithm for UDA training.
@@ III Method
This section details the proposed framework, DEC, grounded in the divide-and-conquer strategy. The framework comprises two components: category models and an ensemble model. Section. 2 discusses the division strategy and category models, which segment images into category masks. Subsequently, Section. III-B explains the ensemble model, designed to fuse predictions derived from category models.
@@ III-A Division Strategy
Employing the divide-and-conquer strategy, we deconstruct the semantic segmentation task into subtasks that group classes into categories and train category models to segment different classes. In particular, we group $N_{C}$ classes into $N_{G}$ categories guided by specific criteria which encompasses various factors, such as how humans semantically label an RGB image by AI tools[20, 21] (e.g. commencing with the background, progressing to large objects, and concluding with smaller ones). To account for non-category classes (classes that do not belong to a particular category), we introduce a class named other category to denote them. It means a specific class is assigned only in one category label for a given pixel, while other category labels maintain the designation as other category. This class helps category models learn features of non-category classes in training, improving the accuracy of classes that belong to a category. Subsequently, we remap source labels $\mathcal{Y^{S}}$ to source category labels $\mathcal{Y}^{S}_{j}$ based on the defined $N_{G}$ categories where the class will be retained in the category label if it belongs to this category. The procedure of generating category labels is shown in Algorithm 1. After generating source category labels, we train target category models, denoted as $\{f_{\theta^{j}}^{T}\}_{j=1}^{N_{G}}$, using existing UDA methods. Each target category model $f_{\theta^{j}}^{T}$ is then employed to segment target images $\mathcal{X}^{T}$ into target category masks $\hat{\mathcal{Y}}^{T}_{j}$, encompassing $N_{j}$ classes.
III-B Ensemble Model
By division strategy detailed in Section. 2, target images $\mathcal{X}^{T}$ are segmented into target category masks $\hat{\mathcal{Y}}^{T}_{j}$. Conflicts arise when different classes are predicted for the same pixel value among the target category masks. Note that popular ensemble methods like majority voting and averaging are not suitable to solve these conflicts. Due to the aforementioned division strategy, category models do not share classes, thus ensemble methods that depends on a consensus between models can not be applied. Hence, we introduce an ensemble model $E_{\theta}$ designed to extract features from each target category mask $\hat{\mathcal{Y}}^{T}_{j}$ and fuse them into the final semantic segmentation mask $\hat{\mathcal{Y}}^{T}$.
@@ III-B Ensemble Model
Training Step. The training step is illustrated in Fig. 2 and implemented in Algorithm 2. In the training dataset ($(\{\hat{\mathcal{Y}}^{S}_{j}\}_{j=1}^{N_{G}},\mathcal{Y}^{S})$, there are $N_{G}$ category masks for one source label. To enable the model to process this kind of input, we stack them (similar to how we treat RGB images) into a pseudo-image $\mathcal{X}^{E}$ where each channel corresponds to a source category mask $\hat{\mathcal{Y}}^{S}_{j}$. The $\mathcal{X}^{E}$ channels follow the same order as the $N_{G}$ categories defined. We use pixel-wise cross-entropy as the loss function $\mathcal{L}^{E}$ to train the ensemble model like semantic segmentation task
@@ IV-B2 Source Category Models
Source category models are employed to train the ensemble model. The network architecture uses DeepLabV3+ [55] with ResNet101 [56] as the backbone, loaded with ImageNet pretrained weights. Training adopts SGD [57] with a learning rate of 2×$\times 10^{-3}$, linear learning rate warmup over 1k iterations, and polynomial decay. The model undergoes training on a batch of eight images with a crop size 1024$\times$512 for 90k iterations.
@@ IV-E1 Source data
TABLE IV: Rare Class Sampling (RCS) and Thing-class ImageNet Feature Distance (FD) are applied to a multi-source dataset, with the results averaged over three random seeds.
@@
TABLE V: The influence of single-source and multi-source training datasets on the ensemble model. We show the performance of GTA5 $\rightarrow$ Cityscapes and Musketeers $\rightarrow$ Cityscapes on the top of MIC.
@@
DEC use multi-source datasets to train category and ensemble models. For category models, as stated in Section I, strategies tailored for single-source synthetic dataset are often susceptible to overfitting and instability. MIC and HRDA exhibit noteworthy advancements in GTA $\rightarrow$ Cityscapes based on RCS and FD while these strategies demonstrate suboptimal performance in multi-source scenarios (see Table IV). On the other hand, the ensemble model needs enough training data since there is no appropriate pre-trained backbone. The variability of single-source dataset is insufficient for developing a robust ensemble model. Table V presents the mIoU of ensemble models trained on single-source and multi-source datasets. The results demonstrate that the ensemble model trained on multi-source datasets effectively fuses category masks into the final mask, e.g., the Human/Cycle category (see Fig. 5 (c)) correctly segments the dotted area, whereas the ensemble model trained with GTA5 fuse dotted area incorrectly (see Fig. 5 (g)). Thus, it is necessary to train the ensemble model with multi-source datasets to obtain robust and reliable final masks.
@@ IV-E2 Division Strategy
TABLE VI: Random division strategy. Each row in the second column is a category.
@@ 4 Rider, Traffic Light, Motorcycle, Car, Truck, Traffic Sign, Pole, Vegetation, other category
DEC groups classes into categories to decrease the complexity of the model and improve segmentation. However, DEC requires a feasible division strategy (including but not limited to Table I) to maintain the contextual information between different classes. Choosing proper categories is not an arbitrary task. We demonstrate in Table VI how randomly pick some classes as a group to generate some categories. It is seen that random division does not push the performance for Musketeers $\rightarrow$ Cityscapes in Table VII (strategy 1-4). Using the four-category division (strategy B+V+H+T), DEC achieves a +1.2 mIoU improvement over MIC. Additionally, we show the performance of other feasible division strategies, such as dividing classes into two categories and three categories in Table VII.
TABLE VII: Performance of selected, random and other feasible division strategies on Musketeers $\rightarrow$ Cityscapes.
Without Division
MIC 97.3 80.4 91.8 57.9 61.1 64.2 68.5 76.6 91.9 52.7 94.8 81.9 59.8 94.8 83.1 90.1 83.0 66.8 76.8 77.5
With Random Division
With B+V+H+T Division
B+V+H+T
97.9
82.2 92.3 59.9 61.5 63.9 69.2 75.0 91.9 53.4 95.0 83.8 67.9 95.5 87.8 90.5 78.5 70.7 78.4 78.7
With Other Feasible Division
These division strategies are generated by combining categories in Table I, e.g., BV denotes a category containing classes from Background and Vehicle; BV+HT means this division strategy splits all classes into two groups. These division strategies achieve a gain of at least $+0.8$ for Musketeers $\rightarrow$ Cityscapes compared to MIC. Fig. 6 provides a qualitative result where these feasible strategies can effectively improve segmentation, e.g., sidewalk.
@@ IV-E3 Ensemble Model
      Backbone       Cityscapes       BDD100K       Mapillary Vista
@@
Backbone: Table VIII illustrates the performance of the ensemble model with various backbones for Musketeers → Cityscapes, BDD100K and Mapillary Vista. ResNet-18 achieves the highest mIoU across all three datasets (79.1, 63.2, and 75.0, respectively). Given its superior performance and fastest training speed among the tested backbones, ResNet-18 is selected as the preferred backbone for the ensemble model.
TABLE IX: The performance of SGD and AdamW optimiser with different learning rates for Musketeers $\rightarrow$ Cityscapes.
@@
Optimizer: Tab. IX shows the SGD and AdamW with different learning rates to train the ensemble model for Musketeers $\rightarrow$ Cityscapes. It can be seen that AdamW is more compatible with the ensemble model than SGD.
TABLE X: The impact of EMA coefficient $\alpha$ for three target domains.
      $\alpha$       Cityscapes       BDD100K       Mapillary Vista
@@
EMA: Tab. X shows the influence of different EMA coefficients $\alpha$ for Musketeers $\rightarrow$ Cityscapes, BDD100K and Mapillary Vista. It is observed that a larger value of $\alpha$ enhances the generalisation of the ensemble model, leading to optimal performance for each target dataset.
IV-F Performance Metrics
TABLE XI: FPS, FLOPs and Parameter size for Input Resolution 1024$\times$512 on RTX 3090
FPS FLOPs Params
(GFLOPs) (M)
HRDA 4.50 410.40 85.69
Category Model (HRDA) 4.50 410.40 85.69
MIC 4.50 410.40 85.69
Category Model (MIC) 4.50 410.40 85.69
Ensemble Model 122.24 37.40 12.30
TABLE XII: Runtime and memory consumption during training and inference on RTX 3090
Training Inference (2048$\times$1024)
Throughput GPU Memory Throughput GPU Memory
(s/iter) (GB) (img/s) (GB)
HRDA 2.55 21.80 1.13 11.93
Category Model (HRDA) 2.35 15.07 1.13 11.93
MIC 3.13 21.73 1.13 11.93
Category Model (MIC) 2.70 15.09 1.13 11.93
Ensemble Model 0.69 17.89 46.30 0.98
Table XI presents the FPS, FLOPs, and parameters for both the category and ensemble models. The category model is built on the existing UDA method, with modifications made to the training labels. Consequently, its FPS, FLOPs, and parameters remain the same as those of the HRDA and MIC models. The ensemble model, with only 12.30M parameters, achieves notably high inference speeds, enabling efficient fusion of the category model outputs. As a result, the overall running speed is primarily dictated by the UDA method used to build the category models—the faster the UDA method, the better our approach performs.
The runtime and memory consumption are presented in Table XII. The category model achieves faster training and reduced memory usage than HRDA and MIC by eliminating unnecessary techniques, such as rare class sampling and the thing-class ImageNet feature distance. Meanwhile, the inference speed and memory consumption remain the same with HRDA and MIC, as no additional parameters are introduced. The ensemble model’s simple architecture completes each training iteration in just 0.69 seconds. Additionally, the ensemble model achieves an inference speed (FPS) of 46.3 for an input resolution of 1024$\times$2048 while consuming only 0.98 GB of memory.
IV-G The Last Mile to SL
@@
Some classes still exhibit gaps compared to SL. For example, the sidewalk, an important class on autonomous driving, shows a 5 points difference from SL on Musketeers $\rightarrow$ Cityscapes. Fig. 7 shows some qualitative results for the sidewalk in both DEC and SL. The error in DEC is primarily found at the bottom of vehicles, where the sidewalk is misclassified as road. We expect these errors to have an insignificant influence on autonomous driving, as the vehicles in these areas are classified correctly. Moreover, since the ground truth is manually labelled, human bias is inevitably introduced into the ground truth. This is why some classes show discrepancies compared to SL, such as terrain, fence, and pole. Fig. 8 illustrates these human biases in Cityscapes.
Furthermore, there are also human bias in Mapillary Vistas. Fig. 9 indicates annotation bias among terrain, building, sky, and vegetation.
These classes are labelled differently between synthetic and real datasets. This explains the large gap (e.g., vegetation, terrain) in our method for Mapillary Vista.